OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = quantity x minus eight divided by quantity x plus seven and g(x) = quantity negative seven x minus eight divided by quantity x minus one.

OpenStudy (mathstudent55):

\(f(x) = \dfrac{x - 8}{x + 7} \) \(g (x) = \dfrac{-7x - 8}{x - 1} \) \(f(g(x)) = \dfrac{\frac{-7x - 8}{x - 1} - 8}{\frac{-7x - 8}{x - 1} + 7} \) \(~~~~~~~~~~~ = \dfrac{(x - 1)(\frac{-7x - 8}{x - 1} - 8)}{(x - 1)(\frac{-7x - 8}{x - 1} + 7)} \) \(~~~~~~~~~~~ = \dfrac{-7x - 8 - 8(x - 1)}{-7x - 8 + 7(x - 1)} \) \(~~~~~~~~~~~ = \dfrac{-7x - 8 - 8x + 8}{-7x - 8 + 7x - 7} \) \(~~~~~~~~~~~ = \dfrac{-15x}{-15} \) \(~~~~~~~~~~~ = x\)

OpenStudy (mathstudent55):

That is how you show f(g(x)) = x. Now do the same for g(f(x)).